OpenStudy (anonymous):

if the point P(-3/5,y) lies on the unit circle P is the second quadrant, what does y equal?

OpenStudy (jdoe0001):

\(\Large \begin{matrix} P\left(-\frac{3}{5}, y\right)\\ P(x\quad, y)\\ P(cosine, sine)\\ P\left(\frac{a}{c},\frac{b}{c}\right)\\ \textit{which means that } -\cfrac{3}{5} \implies \cfrac{a}{c}\\ c^2 = a^2 + b^2 \implies \sqrt{c^2-a^2} = b \end{matrix}\)

OpenStudy (jdoe0001):

keep in mind that when using the pythagorean theorem, you dunno if it's positive or negative, however in this case we know that the point is in the 2nd Quadrant, where "y" is positive